the square o a number, decreased by2 is equal to the negative of that number constuct a quadatic fo that statement and then solve that equation by factoring to find the answer
Answer by jorel1380(3719)

x2-2=-x
x2+x-2=0
(x+2)(x-1)=0
x=-2,1.